OpenStudy (anonymous):

Will give medal and fan you!! :) PLEASE HELP!!! how much distance in (m) will it take a car to stop when a light turns yellow if it is travelling at the speed limit of 60km/h? Assume the driver does not react for 1 seconds in your calculation. In other words, the car continues to travel at the speed limit for 1.0 seconds after the light has turned. Then the driver breaks to stop at an acceleration of 3.0m/s^2

OpenStudy (anonymous):

The distance to cross the intersection is 34m. Assume the length of a typical car is 4.5m

OpenStudy (shamim):

Use the following formula v^2=u^2-2as s=?

OpenStudy (shamim):

Given u=60km/h=?m/s v=0 a=3m/s^2